OpenStudy (anonymous):

plz help me i hve to find circumference of circle n radius is also given 56 plzzzzz help me

OpenStudy (anonymous):

circumference of circle=2 *pi *radius pi=3.14

OpenStudy (anonymous):

circumference of circle=2 *3.14 *56 = ?..

OpenStudy (callisto):

@hubaali Which part you don't understand?

OpenStudy (anonymous):

no i jst wanna ask tht how to do this solution

OpenStudy (anonymous):

sail123 the ans will cme in the decimal

OpenStudy (anonymous):

circumference of circle=2 *22/7 *56 = 44*8

OpenStudy (anonymous):

thn u so mch